OpenStudy (ttop0816):

PLEASE HELPPP (: In triangle ABC, a = 12, B = 25°, and C = 45°. Find b. 5.4 6.6 7.2 22

hartnn (hartnn):

find angle A first A+B+C =180 then you can use sine law

hartnn (hartnn):

\(\large \dfrac{a}{\sin A} = \dfrac{b}{\sin B}\)

OpenStudy (ttop0816):

then would it be A= sin^-1(12) B= sin^-1(25) ??

hartnn (hartnn):

?? first find A! A+B+C =180 B =25 C=45

hartnn (hartnn):

A+B+C = 180 because A,B,C are angles of a triangle.

OpenStudy (anonymous):

looks like you're going in the right direction

OpenStudy (ttop0816):

@hartnn ohhhh!! then A= would be 110! my bad (:

hartnn (hartnn):

yess now apply sine law

hartnn (hartnn):

12/ sin 110 = b/ sin 25 find b

OpenStudy (shinalcantara):



OpenStudy (ttop0816):

sin25*12/sin110=b ?? @hartnn

OpenStudy (anonymous):

im confused now

hartnn (hartnn):

yes, calculate that

hartnn (hartnn):

where r u confused jazi ?

OpenStudy (ttop0816):

@hartnn umm i got this?? 5.396891

hartnn (hartnn):

thats correct :)

hartnn (hartnn):

\(5.396 \approx 5.4\)

OpenStudy (anonymous):

okay, I see now..confused myself guys

OpenStudy (ttop0816):

okay i got it till there! so now how would i find the total answer??

hartnn (hartnn):

just look at the options :P

hartnn (hartnn):

b=5.4!

OpenStudy (ttop0816):

@hartnn hahahah sorry! theres only two more questions that im confused with! could you maybe check them out?? θ is in standard position, and the terminal side passes through the point (3, -5). Find sinθ. -5sqrt31/31 -5sqrt34/34 -3/5 -4/5

OpenStudy (ttop0816):

find sintheta!

hartnn (hartnn):

point = (r cos theta , r sin theta) so to get sin theta, do -5/r know how to find 'r' ?

OpenStudy (ttop0816):

sin^-1(-5)??

hartnn (hartnn):

\(r = \sqrt {3^2 +(-5)^2}\)

OpenStudy (ttop0816):

i think i got the answer!! -5sqrt34/34

hartnn (hartnn):

and thats correct :)